# Language-guided Open-world Video Anomaly Detection

### v1

@@ 4 Method: LaGoVAD
We implement the language-guided VAD paradigm via LaGoVAD. We first introduce the overall architecture, followed by details of two proposed regularization terms (i.e., dynamic video synthesis and negative mining). As illustrated in Fig. 3, we take video $v$ and anomaly definition $z$ as inputs. The video is synthesized by a non-parametric dynamic video synthesis module. The anomaly definition is a category set $z=\{z_{0},z_{1},\dots,z_{C-1}\}$, where each class $z_{i}$ is defined by a class name or a description and $C$ is the number of categories in a certain definition. During training, we randomly choose either the class names or the anomaly descriptions within a batch as the definition. Specially, $z_{0}$ denotes the normal class. We extract and encode features of videos with $\mathcal{F}$, which includes a pretrained CLIP image encoder [22] and a Transformer-based temporal encoder. And the text features are extracted with CLIP text encoder $\mathcal{G}$. Then, the encoded features are fused by a Transformer-based fusion module $\mathcal{U}$. Finally, the fused features are fed into a detection head $\mathcal{H}^{\text{bin}}$ to obtain the anomaly score $y^{\text{bin}}\in\mathbb{R}^{L\times 1}$ and a classification head $\mathcal{H}^{\text{mul}}$ to obtain the classification probability $y^{\text{mul}}\in\mathbb{R}^{L\times C}$, where $L$ is the length of video. Formally,
$[v,y^{\text{p}}]=\text{Synthesis}(N,A),$ (5)
$v^{t}=\mathcal{F}(v),\quad z^{t}=\mathcal{G}(z),$ (6)
$[v^{u},z^{u}]=\mathcal{U}(v^{t},z^{t}),$ (7)
$y^{\text{bin}}=\mathcal{H}^{\text{bin}}(v^{u}),\quad y^{\text{mul}}=\mathcal{H}^{\text{mul}}(v^{u},z^{u})$ (8)
where $N,A$ are normal and abnormal video sets, $\text{Synthesis}(\cdot,\cdot)$ is the dynamic video synthesis module, $y^{\text{p}}$ is the pseudo-label generated during synthesis, $v^{t},z^{t}$ are encoded features and $v^{u},z^{u}$ are fused features. In this paper, we address the challenge of concept drift in open-world scenarios to enhance model generalization. To this end, we build a model with intentionally simplified architecture.
During training, we optimize the model through four losses under weak supervision. Following [21, 34, 35, 25, 9], we use multiple instance learning loss $\mathcal{L}_{\text{MIL}}$ to optimize detection. And we employ MIL-align loss $\mathcal{L}_{\text{MIL-align}}$ to optimize classification following [35, 34]. Our paradigm operates in multimodal joint spaces ($P(v,z,y)$) that inherently suffer from exponentially decaying sample density, thereby inducing overfitting problems. Specifically, the algorithm may establish a wrong mapping or suppress a certain modality. Therefore, we leverage more diverse videos via a dynamic video synthesis loss $\mathcal{L}_{\text{dvs}}$ to learn better mappings. We also incorporate a contrastive learning loss with hard negative mining $\mathcal{L}_{\text{neg}}$ to better align two modalities and mitigate the imbalance of two modalities. Formally,
@@ 4.1 Dynamic Video Synthesis
In real-world scenarios, anomalies typically occupy only a small portion of a lengthy video, whereas current datasets predominantly contain videos with high anomaly ratios due to web-sourced data limitations. To mitigate this bias, we dynamically synthesize videos with varying durations and compute a loss based on the pseudo label generated during synthesis, which is inspired by Mosaic [3] and Cutout [11] augmentation. The module initially determines whether to generate a normal or abnormal video, followed by specifying the number of segments. It then selects an anchor video and randomly selects similar videos from k-nearest neighbors to construct a semantically coherent sequence, where the anchor’s position is transformed to a binary pseudo label $y^{\text{p}}\in{\{0,1\}}^{L}$, where $L$ denotes the feature length. Notably, the distance metrics required for retrieval are pre-computed, effectively reducing computational overhead during training. Finally, a dynamic video synthesis is calculated as:
@@ 4.2 Contrastive Loss with Hard Negative Mining
$\tilde{v}^{\text{pos}}=\sum_{i}^{L}v^{t}_{i}\frac{\exp{(y^{\text{bin}}_{i}/\eta)}}{\sum_{j}^{L}\exp{(y^{\text{bin}}_{j}/\eta)}},$ (13)
$\tilde{v}^{\text{neg}}=\sum_{i}^{L}v^{t}_{i}\frac{\exp{(-y^{\text{bin}}_{i}/\eta)}}{\sum_{j}^{L}\exp{(-y^{\text{bin}}_{j}/\eta)}},$ (14)
where $v^{t}_{i}$ denotes the $i$-th feature in $v^{t}$, $\eta$ denotes the temperature, $\tilde{v}^{\text{pos}}$ denotes the aggregated foreground feature and $\tilde{v}^{\text{neg}}$ denotes the aggregated background feature. The background feature in an abnormal video is the normal part of it, which could be considered as the hard negative to its corresponding anomaly description. Therefore, we obtain $\tilde{v}^{\text{pos}}$ of all samples and $\tilde{v}^{\text{neg}}$ of only abnormal samples in a batch, forming $\tilde{V}\in\mathbb{R}^{(B_{1}+B_{2})\times E}$, where $B_{1}$ is the batch size, $B_{2}$ is the number of abnormal videos in a batch, and $E$ is the feature dimension. We also obtain text features before fusing, forming $\tilde{Z}\in\mathbb{R}^{B_{2}\times E}$. The contrastive loss is as follows:
$S=\mathop{Norm}(\tilde{V})\times\mathop{Norm}(\tilde{Z}),$ (15)
$\mathcal{L}_{t\rightarrow v}=-\sum^{B_{2}}_{i}\log\frac{\exp{(S_{i,i}/\tau)}}{\sum^{B_{1}+B_{2}}_{j}\exp{(S_{j,i}/\tau)}},$ (16)
$\mathcal{L}_{v\rightarrow t}=-\sum^{B_{2}}_{j}\log\frac{\exp{(S_{j,j}/\tau)}}{\sum^{B_{2}}_{i}\exp{(S_{i,j}/\tau)}},$ (17)
@@
where $\mathop{Norm}$ is L2 normalization and $\tau$ denotes the temperature.
During inference, the user can input either descriptions or class names as the anomaly definition. For the classification head, we select the minimum value of the normal class and the maximum value of the abnormal class over the temporal axis and use these values after applying Softmax as probabilities. More details are provided in supp(Sec. B).
@@ 5 Dataset: PreVAD
As in Eq. 4, our new paradigm requires diverse $(v,z,y)$ triples for training. Therefore, we propose PreVAD—a large-scale video anomaly detection dataset, which is collected through a scalable curation pipeline.
@@ 5.1 Data Curation Pipeline
Our pipeline encompasses three stages: source, cleansing, and annotation, leveraging multiple foundation models to ensure cost efficiency while maximizing data quality. We aggregate videos from three sources. We first leverage existing video-text datasets [36, 29, 16, 44] to retrieve anomaly videos through text-based video retrieval. Second, we expand the collection through curated web resources, including 1) accident compilations, self-defense tutorials, and fail videos; 2) driving vlogs and travel documentaries; 3) violence recognition datasets [7]. Last, we obtain normal surveillance videos from YouTube streams and traffic camera videos from government-released road camera streams.
In the cleansing stage, we first use PySceneDetect to remove irrelevant segments such as intros and outros. Next, a multimodal LLM (MLLM) generates detailed video descriptions, and a vision-language model (VLM) verifies the consistency between the descriptions and video content. Finally, an LLM evaluates the descriptions to confirm the presence of anomalies, decreasing false positives and ensuring high-quality data.
The annotation stage involves a weakly supervised labeling approach with hybrid human-AI annotation. We annotate each video with a video-level category and annotate each validation-set video with a frame-level label. Additionally, an MLLM generates fine-grained anomaly descriptions for each video under constrained prompts, which include human-labeled categories, and is able to guide MLLM to describe only the anomaly in the video. Notably, we do not additionally label a test set, as we will conduct zero-shot evaluations on other existing VAD datasets. More details can be found in the supp (Sec. C)
Figure 4: The statistics of PreVAD.
Figure 5: Comparison between PreVAD and existing datasets.
Figure 6: A sample of PreVAD, which includes multi-level category label and precise description of the anomaly.
@@ 5.2 Dataset Statistics
Our model stands out on a larger scale, with a wider variety of anomaly types and high-quality anomaly descriptions.
Scale. As in Fig. 4b, PreVAD comprises 35,279 videos, with 11,979 abnormal videos and 23,300 normal videos, partitioned into training and validation sets, which is the largest video anomaly dataset up to now.
Anomaly Types. The diversity of a dataset is related to both the granularity and breadth of its taxonomy. Our dataset features a hierarchical taxonomy comprising 7 first-level categories (i.e., Violence, Vehicle Accident, Fire-related Accident, Robbery, Daily Accident, Animal-related Violence, Production Accident) and 35 subcategories (e.g., carjacking, mugging, sport fail, war). Our taxonomy spans minor (e.g., fall to the ground) to severe anomalies (e.g., shooting), covering most of the common types of anomaly under surveillance. Moreover, the hierarchical design helps model learning diverse definitions.
Anomaly Descriptions. Each abnormal video is annotated with a text description, which has a total vocabulary size of 5,298 words and an average of 22.9 words per description. As shown in Figure 6, our annotation accurately describes the abnormal objects and behaviors in a fine-grained manner, which enables the model to learn better alignment.
Statistics. The videos have an average duration of 21.38 seconds (as in Fig. 4a), with the majority of them clustering within 5 to 20 seconds. And there are over 2000 videos that exceed one minute. Unlike UCF-Crime [25] and XD-Violence [32], PreVAD does not include untrimmed videos that last for tens of minutes or even several hours, as this would lack diversity and be detrimental to both training and validation. As shown in Fig. 4c, most of the videos are from existing video-text datasets or streaming, significantly reducing the overhead of manual clipping and retrieval. Unlike datasets assembled by merging existing VAD datasets, PreVAD obtains videos independently, enabling cross-dataset validation as a new generalization benchmark. Please refer to the supplementary material for more details (Sec. C).
@@ 6 Experiments
Table 2: Zero-shot comparison with other methods in video anomaly detection. We reproduced models marked with $\star$ using their open-source codes, and the rest of the data are taken from their publications. Results marked with $\dagger$ are from [40].
Methods Training-set Test-set
UCF XD MSAD UBN DoTA TAD LAD
OVVAD[34] AIGC+XD 82.42 - - - - - -
LaGoVAD PreVAD+XD 82.81 - - - - - -
OVVAD AIGC+UCF - 63.74 - - - - -
LaGoVAD PreVAD+UCF - 76.28 - - - - -
CLIP[22] - 53.16 17.83 - - - - -
LLaVA1.5[15] - 72.84 50.26 - - - - -
LAVAD[40] - 80.28 62.01 - - - - -
MIL[25] UCF - - 49.50 - - -
CMRL[8] UCF - 46.74 - - - - -
MultiDomain[9] Multiple 78.55 - - - - 79.21 77.36
PEL[21] UCF - 43.53 79.82 54.02 53.05 86.27 69.99
PEL XD 54.52 - 68.25 49.55 44.97 43.02 30.82
VadCLIP[35] UCF - 58.29 88.09 56.24 50.93 74.46 74.29
VadCLIP XD 80.16 - 88.48 57.41 49.00 83.56 74.46
VadCLIP PreVAD 79.37 67.43 89.79 55.66 50.59 85.96 75.02
LaGoVAD PreVAD 81.12 74.25 90.41 58.07 62.60 89.56 78.91
@@ 6.1 Experiment Setup
Datasets & Metrics
This study aims to enhance the open-world generalization of VAD models. We therefore conduct comprehensive evaluations across seven datasets: UCF-Crime (UCF) [25], XD-Violence (XD) [32], MSAD [45], UBNormal (UBN) [1], DoTA [38], TAD [17], and LAD [27], which encompass diverse anomaly types. The validation set of our proposed PreVAD is utilized for in-domain performance analysis and ablation studies. During zero-shot evaluation, we use manual designed prompts based on the class name of the corresponding dataset as the anomaly definition. For detection metrics, we follow others using Average Precision (AP) for XD-Violence, while employing Area Under the Curve of the frame-level receiver operating characteristic (AUC) for other datasets. For classification metrics, there is currently no consensus evaluating classification performance in video anomaly detection. [35] uses mAP, and [34] uses accuracy for abnormal videos, but they neglect the evaluation of classifying normal videos. Therefore, we measure using accuracy and F1-score on both abnormal and normal videos.
Baselines
Given the absence of prior zero-shot evaluations on all seven datasets, we reproduce PEL [21] and VadCLIP [35] using publicly available weights and code for comparison. For classification, we further benchmark against zero-shot action recognition models [28, 23]. To ensure fairness in comparisons with CLIP-based methods, we use the identical ViT/B-16 variant.
Details of the implementation, evaluation datasets, and reproduced baselines are provided in the supp (Secs. B.2, D and E)
Table 3: Zero-shot comparison with other methods in video anomaly classification. The results for other models are obtained with their open-source codes and weights. $\dagger$ denotes providing ground-truth segments when classifying abnormal videos.
@@
CLIP [22] - 19.31 12.08 56.25 45.04
CLIP - 20.34 11.10 64.25 54.61
ActionCLIP [28] K400 18.62 16.12 38.75 37.11
ActionCLIP K400 19.31 13.85 41.37 38.58
ViFi-CLIP [23] K400 20.34 15.67 53.75 50.33
VadCLIP [35] UCF - - 46.38 26.16

### v2

@@ 4 Method: LaGoVAD
We implement the language-guided VAD paradigm (Eq. 4) via LaGoVAD. We first introduce the overall architecture, followed by details of two proposed regularization terms.
As illustrated in Fig. 2, we take video $v$ and anomaly definition $z$ as inputs. The video is synthesized by a non-parametric dynamic video synthesis module. The anomaly definition is a category set $z=\{z_{0},z_{1},\dots,z_{C-1}\}$, where each class $z_{i}$ is defined by a class name or a description and $C$ is the number of categories in a certain definition. During training, we randomly choose either the class names or the anomaly descriptions within a batch as the definition. We extract and encode features of videos with $\mathcal{F}$, which includes a pretrained CLIP image encoder (Radford et al., 2021) and a Transformer-based temporal encoder. And the text features are extracted with CLIP text encoder $\mathcal{G}$. Then, the encoded features are fused by a Transformer-based fusion module $\mathcal{U}$. Finally, the fused features are fed into a binary detection head $\mathcal{H}^{\text{bin}}$ to obtain the anomaly score $y^{\text{bin}}\in\mathbb{R}^{L\times 1}$ and a multi-classification head $\mathcal{H}^{\text{mul}}$ to obtain the classification probability $y^{\text{mul}}\in\mathbb{R}^{L\times C}$, where $L$ is the length of video. Formally,
$v,y^{\text{p}}]=\text{Synthesis}(N,A),$ (5)
$v^{t}=\mathcal{F}(v),\quad z^{t}=\mathcal{G}(z),\quad[v^{u},z^{u}]=\mathcal{U}(v^{t},z^{t}),$ (6)
$y^{\text{bin}}=\mathcal{H}^{\text{bin}}(v^{u}),\quad y^{\text{mul}}=\mathcal{H}^{\text{mul}}(v^{u},z^{u}),$ (7)
where $N,A$ are normal and abnormal video sets, $\text{Synthesis}(\cdot,\cdot)$ is the dynamic video synthesis module, $y^{\text{p}}$ is the pseudo-label generated during synthesis, $v^{t},z^{t}$ are encoded features and $v^{u},z^{u}$ are fused features. This work prioritizes addressing the challenge of concept drift over designing complex architectures. Consequently, we adopt a simplified network.
During training, we optimize the model through four losses under weak supervision. Following (Wu et al., 2024b; c), we calculate multiple instance learning loss $\mathcal{L}_{\text{MIL}}$ (with $y^{\text{bin}}$) and MIL-align loss $\mathcal{L}_{\text{MIL-align}}$ (with $y^{\text{mul}}$) to optimize temporal binary detection and video-level multi-class classification. Our paradigm operates in multimodal joint spaces $P(V,Z,Y)$ that inherently suffer from exponentially decaying sample density, thereby inducing overfitting problems. Specifically, the algorithm may establish a wrong mapping or suppress a certain modality. Therefore, we leverage more diverse videos via a dynamic video synthesis loss $\mathcal{L}_{\text{dvs}}$ to learn better mappings. We also incorporate a contrastive learning loss with hard negative mining $\mathcal{L}_{\text{neg}}$ for better alignment. Formally,
@@ 4.1 Dynamic Video Synthesis
In real-world scenarios, anomalies typically occupy only a small portion of a lengthy video, whereas current datasets predominantly contain videos with high anomaly ratios due to web-sourced data limitations. To mitigate this bias, we dynamically synthesize videos with varying durations and compute a loss based on the pseudo label generated during synthesis. The module initially determines whether to generate a normal or abnormal video, followed by specifying the number of segments. It then selects an anchor video and randomly selects similar videos from k-nearest neighbors to construct a semantically coherent sequence, where the anchor’s position is transformed to a binary pseudo label $y^{\text{p}}\in{\{0,1\}}^{L}$, where $L$ denotes the feature length. Notably, the distance metrics required for retrieval are pre-computed, effectively reducing computational overhead during training. Finally, a dynamic video synthesis is calculated as:
$\mathcal{L}_{\text{dvs}}=$ $-\hat{y}\log\textstyle\sum_{i\in\Omega^{a}_{k}}\sigma(y^{\text{bin}}_{i})/k-(1-\hat{y})\log(1-\textstyle\sum_{i\in\Omega^{n}_{k}}\sigma(y^{\text{bin}}_{i})/k)$ (9)
@@ 4.2 Contrastive Loss with Hard Negative Mining
$\tilde{v}^{\text{pos}}=\sum_{i}^{L}v^{t}_{i}\frac{\exp{(y^{\text{bin}}_{i}/\eta)}}{\sum_{j}^{L}\exp{(y^{\text{bin}}_{j}/\eta)}},\quad\tilde{v}^{\text{neg}}=\sum_{i}^{L}v^{t}_{i}\frac{\exp{(-y^{\text{bin}}_{i}/\eta)}}{\sum_{j}^{L}\exp{(-y^{\text{bin}}_{j}/\eta)}},$ (11)
where $v^{t}_{i}$ denotes the $i$-th feature in $v^{t}$, $\eta$ denotes the temperature, $\tilde{v}^{\text{pos}},\tilde{v}^{\text{neg}}$ denote the aggregated foreground/background feature. The background feature in an abnormal video is the normal part of it, which could be considered as the hard negative to its corresponding anomaly description. Therefore, we obtain $\tilde{v}^{\text{pos}}$ of all samples and $\tilde{v}^{\text{neg}}$ of only abnormal samples in a batch, forming $\tilde{V}\in\mathbb{R}^{(B_{1}+B_{2})\times E}$, where $B_{1}$ is the batch size, $B_{2}$ is the number of abnormal videos in a batch, and $E$ is the feature dimension. We also obtain the text features before fusing, forming $\tilde{Z}\in\mathbb{R}^{B_{2}\times E}$. The contrastive loss is as follows:
$\mathcal{L}_{t\rightarrow v}=-\sum^{B_{2}}_{i}\log\frac{\exp{(S_{i,i}/\tau)}}{\sum^{B_{1}+B_{2}}_{j}\exp{(S_{j,i}/\tau)}},\quad\mathcal{L}_{v\rightarrow t}=-\sum^{B_{2}}_{j}\log\frac{\exp{(S_{j,j}/\tau)}}{\sum^{B_{2}}_{i}\exp{(S_{i,j}/\tau)}},$ (12)
@@
[…]
During inference, the user can input either descriptions or class names as the anomaly definition. For the classification head, we select the minimum value of the normal class and the maximum value of the abnormal class over the temporal axis and use these values after applying Softmax as probabilities. More architecture details are provided in supp (Sec. B.1).
@@ 5.1 Data Curation Pipeline
We propose PreVAD—a large-scale pretraining VAD dataset to provide diverse $(v,z,y)$ triples for training, which is collected through a scalable curation pipeline. The proposed pipeline encompasses three stages: source, cleansing, and annotation.
We aggregate videos from three sources: First, we retrieve anomaly videos from existing large-scale video-text datasets (Xu et al., 2016; Wang et al., 2019; Liu et al., 2025; Zhu et al., 2024a) utilizing their text annotation. Second, we expand the collection through curated web resources, including 1) accident compilations and fail videos; 2) driving and travel vlogs; 3) violence recognition videos. Last, we obtain normal surveillance videos from YouTube streams and traffic camera streams.
In the cleansing stage, we first remove irrelevant segments such as intros and outros with automatic tools. Next, a multimodal LLM (MLLM) generates detailed video descriptions, and a vision-language model verifies the consistency between the descriptions and video. Finally, an LLM evaluates the descriptions to confirm the presence of anomalies, decreasing hallucinations.
In the annotation stage, we employ a hybrid human-AI approach. First, we annotate each video with a category label. Then, using this label as a constraint, we prompt an MLLM to generate fine-grained descriptions of the anomalies, ensuring focused and relevant output. We also conduct frame-level annotations for the validation set. Notably, we do not additionally label a test set, as we will conduct zero-shot evaluations on other existing VAD datasets. More details can be found in the supp Sec. C.
Figure 3: The statistics, comparisons and a data sample of the proposed PreVAD.
@@ 5.2 Dataset Statistics
Our dataset stands out for its large scale, wide variety of anomalies, and high-quality descriptions.
Scale. PreVAD comprises 35,279 videos, spanning a total duration of 209.5 hours, with 11,979 abnormal videos and 23,300 normal videos, partitioned as shown in Fig. 3a, which is the largest video anomaly dataset up to now, as compared in Fig. 3b.
Anomaly Types. Our dataset’s diversity stems from a hierarchical taxonomy with 7 first-level categories (i.e., Violence, Vehicle Accident, Fire-related Accident, Robbery, Daily Accident, Animal-related Violence, Production Accident) and 35 subcategories (e.g., carjacking, mugging, sport fail, war), spanning from minor (e.g., fall to the ground) to severe (e.g., shooting) anomalies, which covers a broad range of scenarios.
Anomaly Descriptions. Each abnormal video is annotated with a text description, which has a total vocabulary size of 5,298 words and an average of 22.9 words per description. As shown in Fig. 3c, our annotation accurately describes the abnormal objects and behaviors in a fine-grained manner.
Sources. As shown in Fig. 3d, most of the videos are from existing video-text datasets or streaming, significantly reducing the overhead of manual clipping and retrieval. PreVAD also obtains videos independently without merging existing VAD datasets, enabling cross-dataset validation as a new generalization benchmark. We provide more details of PreVAD in the supp Sec. C.
@@ 6 Experiments
Table 2: Comparison in temporal binary anomaly detection under Protocol 1. Results marked with $\dagger$ are taken from their publications and results marked with $\ddagger$ are from Zanella et al. (2024).
Refer to caption Methods Training-set Test-set UCF XD MSAD UBN DoTA TAD LAD (AUC) (AP) (AUC) (AUC) (AUC) (AUC) (AUC) OVVAD AIGC+XD 82.42 - - - - - - LaGoVAD PreVAD+XD 82.81 - - - - - - OVVAD AIGC+UCF - 63.74 - - - - - LaGoVAD PreVAD+UCF - 76.28 - - - - - CLIP - 53.16 17.83 - - - - - LLaVA1.5 - 72.84 50.26 - - - - - LAVAD - 80.28 62.01 - - - - - CMRL UCF - 46.74 - - - - - MultiDomain Multiple 78.55 - - - - 79.21 77.36 PEL UCF - 43.53 79.82 54.02 53.05 86.27 69.99 PEL XD 54.52 - 68.25 49.55 44.97 43.02 30.82 VadCLIP UCF - 58.29 88.09 56.24 50.93 74.46 74.29 VadCLIP XD 80.16 - 88.48 57.41 49.00 83.56 74.46 VadCLIP PreVAD 79.37 67.43 89.79 55.66 50.59 85.96 75.02 LaGoVAD PreVAD 81.12 74.25 90.41 58.07 62.60 89.56 78.91
@@ 6.1 Experiment Setup
Datasets
We conduct comprehensive zero-shot evaluations across seven datasets: UCF-Crime (UCF) (Sultani et al., 2018), XD-Violence (XD) (Wu et al., 2020), MSAD (Zhu et al., 2024b), UBNormal (UBN) (Acsintoae et al., 2022), DoTA (Yao et al., 2022), TAD (Lv et al., 2021), and LAD (Wan et al., 2021), which encompass diverse anomaly types. The validation set of our proposed PreVAD is also utilized for in-domain analysis and ablations. More details are in supp Sec. D.1.
Evaluation Protocols
We evaluate the open-world capability with two zero-shot protocals: Protocol 1: Testing on multiple test sets separately, each representing a distinct scenario (e.g., TAD for traffic scenarios), which evaluates the overall performance under concept drift, unseen categories, feature distribution shifts, etc. Protocol 2: Testing on a dataset with varying anomaly definitions, where in each definition a different subset of anomaly categories is redefined as normal. The final performance is averaged across five such definitions (denoted as drift@5), specifically evaluating the model’s robustness to concept drift. The differences of test sets and the selected subsets are detailed in supp Sec. D.2. During evaluation, we use manual designed prompts based on the class name of the corresponding dataset as the anomaly definition.
Comparative Methods
For Protocol 1, we compare against traditional methods (PEL (Pu et al., 2024), VadCLIP (Wu et al., 2024c)), along with scene-dependent (CMRL (Cho et al., 2023)), zero-shot (LAVAD (Zanella et al., 2024)), open-vocabulary (OVVAD (Wu et al., 2024b)), and multi-domain generalization (MultiDomain (Cho et al., 2024)) approaches. We also include open-vocabulary action recognition methods (ActionCLIP (Wang et al., 2021), ViFi-CLIP Rasheed et al. (2023)) for multi-class comparison. For Protocol 2, as most methods do not support user-provided anomaly definition, comparisons are primarily made with LLM-based (Qwen2-VL (Wang et al., 2024a), Qwen2.5-VL (Bai et al., 2025), LAVAD, HolmesVAU Zhang et al. (2025)) and multi-modal methods (VadCLIP). All results are based on their open-source codes and weights, detailed in Sec. E.
Metrics
For binary detection metrics and without additional annotations, we follow existing works using Average Precision (AP) for XD-Violence and using Area Under the Curve of the frame-level receiver operating characteristic (AUC) for other datasets. For multi-class classification metrics, we use multi-class accuracy and F1-score on both abnormal and normal videos.
Details of implementation of our method are in supp Sec. B.2.
Table 3: Comparison in video-level multi-class classification on UCF-Crime and XD-Violence under Protocol 1.
@@
CLIP - 19.31 12.08 56.25 45.04
ActionCLIP K400 18.62 16.12 38.75 37.11
ViFi-CLIP K400 20.34 15.67 53.75 50.33
VadCLIP UCF - - 46.38 26.16
@@
Table 4: Comparison in temporal binary anomaly detection on XD and MSAD under Protocol 2, specifically evaluating robustenss to concept drift.
Method XD-drift@5 MSAD-drift@5
AUC AP AUC AP
Qwen2-VL-7B 0.604 0.175 0.654 0.229
Qwen2.5-VL-7B 0.627 0.206 0.631 0.224
VadCLIP 0.813 0.358 0.852 0.188
HolmesVAU - - 0.843 0.343
LAVAD 0.817 0.348 0.722 0.315
LaGoVAD 0.857 0.371 0.856 0.401